OpenStudy (anonymous):

In a for the purpose of this question irrelevant physics exercise I end up needing to derive the following function to t: -L_1*sin(theta)*w-(K_1*sin(2*theta)*D*w)/(2*(L-D*cos(theta))^2) with L_1= K_1/(L-d*cos(theta)) L and D are known constants w=d(theta)/dt, a=dw/dt Now, I know how I could solve this using the standard chainrule, but on an exam that would a) take way too long (strong time restrictions), b) lead to high level of possible little errors. Is there way, maybe with trigfunctions or something else, in which I could solve this derivative more quickly?

OpenStudy (anonymous):

K_1 is also a known constant

OpenStudy (anonymous):

\[-L_1\cdot\sin(\theta)\cdot w-\frac{K_1\cdot\sin(2\theta)\cdot D\cdot w}{2\cdot(L-D\cdot\cos(\theta))^2}\] with \(L_1= \dfrac{K_1}{L-d\cdot\cos(\theta))}\) L and D are known constants\[w=\frac{\textrm{d}\theta}{\textrm{d}t}, a=\frac{\textrm{d}w}{\textrm{d}t}\]

OpenStudy (anonymous):

How do you type tex on this site btw?

OpenStudy (anonymous):

Sorry, I can't help, but I thought others may find this a little easier to read.

OpenStudy (anonymous):

(I just meant which tags to use, like [math][/math] or $$)

OpenStudy (anonymous):

Oh \(\TeX\) :). For in-line typing, encapsulate \(\LaTeX\) code in `\(...\)`. For block style typing, encapsulate in `\[...\]`. There are many restrictions, though. You can't use packages, and some of the language has different idioms than standard \(\LaTeX\). You can still use environments and minipages and such. Check out http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A) for mor info.

OpenStudy (anonymous):

Thanks. Now, the actual question though :p

OpenStudy (anonymous):

Yea that's what I was referring when I said I can't help :(

OpenStudy (anonymous):

k, thanks anyway